# Constraints on the $\gamma$-parameter for the vacuum solution of Cotton gravity with geodesics and shadows

### v1

@@ VII Theoretical Lyapunov Exponent
In this section, we explore the Lyapunov exponent of nearly bound orbits, a quantity that measures the instability scale of such orbits. Since these instabilities take place near the photon sphere, the Lyapunov exponent tracks a local effect and is therefore independent of the observer’s distance to the black hole. We begin our analysis by considering the effect of a small radial perturbation near the photon sphere (i.e. $r=\delta r+r_{p}\ ,\delta r\ll r_{p}$) and apply it to Eq. (14). We perform a second-order Taylor series expansion of the effective potential $V_{\rm eff}$ around $r_{p}$
$V_{\rm eff}(r)\simeq V_{\rm eff}(r_{p})+V^{\prime}_{\rm eff}(r_{p})\delta r+\frac{1}{2}V^{\prime\prime}_{\rm eff}(r_{p})(\delta r)^{2},$ (55)
where the first term of the expansion corresponds to the critical impact parameter $b_{c}$ and the second term disappears, since unstable circular null geodesics correspond to critical maxima of the effective potential. The prime notation above denotes a derivative with respect to the radial coordinate. Substituting this result into Eq. (14), leads to the following expression
$\left(\frac{d\delta r}{d\phi}\right)^{2}=-\frac{1}{2}C^{2}(r_{p})V^{\prime\prime}_{\rm eff}(r_{p})(\delta r)^{2}.$ (56)
@@
$\pi\frac{d\delta r}{d\phi}=\lambda\delta r,$ (57)
whose solutions are of an exponential form, $\delta r_{n}=\delta r_{0}e^{\lambda n}$, we obtain the Lyapunov exponent $\lambda$ which quantifies the instability scale of a small perturbation. Computing the second derivative of the effective potential and utilizing the critical curve equation $C^{\prime}(r_{p})A(r_{p})-C(r_{p})A^{\prime}(r_{p})=0$ [59], yields the following expression for the theoretical Lyapunov exponent
$\lambda=\pi\left(\frac{AC^{\prime\prime}-A^{\prime\prime}C}{2}\right)^{\frac{1}{2}}.$ (58)
For the Schwarzschild geometry $\gamma=\pi$, indicating that an initial perturbation grows (in radial distance) by a factor $e^{\gamma}$ with each half-orbit around the black hole. For the Cotton geometry, the Lyapunov exponent reads
$\lambda=\pi(1+\gamma r_{p})^{\frac{1}{2}}.$ (59)
This result supports the qualitative analysis done in the previous sections. The $\lambda$ exponent of the CG black hole, differs from the Schwarzschild black hole by an order of magnitude proportional to $\gamma$. Consequently, despite the exponential behaviour driven by the Lyapunov exponent, the deviations between these two geometries are captured by photon rings of a similarly order, requiring many half-orbits around the black hole before becoming significantly different. The gravitational lensing causing the $n=1$ and $n=2$ photon rings is far from being sensitive to the contribution of $\gamma$. Thus, one cannot hope to be able to distinguish between the Schwarzschild and CG geometries via their $n=1$ and $n=2$ ring features (i.e. relative intensities, relative widths), which are the ones that could be within reach of future interferometric detectors [60].
@@ VIII Conclusion
To study the impact of this result in geodesic effects, we first computed the exact expression for the deflection angle in the form of elliptic integrals. There are two important considerations about this result. Firstly, in order to obtain the result of the Schwarzschild solution, one must first assume $\gamma=0$ and additionally $u_{R}$ and $u_{S}=0$ (which corresponds to the consideration of infinite distances). This is because the result for Schwarzschild is usually determined using the definition for infinite distances, which is not possible in the case of Cotton gravity given its non-asymptotically flat character. Secondly, the deflection angle can be written as a function of the impact parameter $b$, though under an explicit cumbersome form. Furthermore, we computed the deflection angle using a Taylor series approximation to solve the integrals, which yields a result which is valid for the weak field regime where $M/b\ll 1$. We have considered a linear approximation for the parameter $\gamma$ in addition to the series with respect to the mass (we expect that $\gamma$ has in reality a much smaller order of magnitude than mass and distances), and checked the consistence of the analytical approximation with the numerical integration of the deflection angle. Overall, we find that the deflection angle remains unchanged for values of the CG parameter $\gamma<10^{-14}$.
@@
In looking for further settings to search for the presence of any differences, we studied the isoradial curves of a CG black hole against those of a Schwarzschild black hole, utilizing a generalization of the Poutanen formula. We found that there is a full overlap of the isoradial curves between the CG and Schwarzschild geometries, confirming the absence of any visual differences in this regard. We also considered the theoretical Lyapunov exponent for a Cotton gravity black hole, associated to the instability scale of photon rings. We derived its exact expression by perturbing the photon equation of motion about the photon sphere radius. The CG theoretical Lyapunov exponent was found to differ from Schwarzschild’s by a factor proportional to CG’s $\gamma$ parameter, and thus photon rings do not help in discriminating between both geometries, at least for those rings ($n=1$ and $n=2$) that could be expected to be measured in the last decades via interferometric means.
@@ Acknowledgements.
MER thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil, for partial financial support. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. FSNL acknowledges support from the Fundação para a Ciência e a Tecnologia (FCT) Scientific Employment Stimulus contract with reference CEECINST/00032/2018, and funding through the research grants UIDB/04434/2020, UIDP/04434/2020 and PTDC/FIS-AST/0054/2021. DRG is supported by the Spanish Agencia Estatal de Investigación Grant No. PID2022-138607NB-I00, funded by MCIN/AEI/10.13039/501100011033, UE (European Union), ERDF (European Regional Development Fund) A way of making Europe.

### v2

@@ VII Theoretical Lyapunov Exponent
In this section we explore the Lyapunov exponent of nearly bound orbits, a quantity that measures the instability scale of such orbits. Since these instabilities take place near the photon sphere, the Lyapunov exponent tracks a local effect and is therefore independent of the observer’s distance to the black hole. We begin our analysis by considering the effect of a small radial perturbation near the photon sphere (i.e. $r=\delta r+r_{p}\ ,\delta r\ll r_{p}$) and apply it to Eq.(14). We perform a second-order Taylor series expansion of the effective potential $V_{eff}$ around $r_{p}$
$V_{eff}(r)\simeq V_{eff}(r_{p})+V^{\prime}_{eff}(r_{p})\delta r+\frac{1}{2}V^{\prime\prime}_{eff}(r_{p})(\delta r)^{2},$ (50)
where the first term of the expansion corresponds to the critical impact parameter $b_{c}$ and the second term disappears, since unstable circular null geodesics correspond to critical maxima of the effective potential. The prime notation above denotes a derivative with respect to the radial coordinate. Substituting this result into Eq.(14), leading to the following expression
$\left(\frac{d\delta r}{d\phi}\right)^{2}=-\frac{1}{2}\frac{C^{2}(r_{p})}{A(r_{p})B(r_{p})}V^{\prime\prime}_{eff}(r_{p})(\delta r)^{2}.$ (51)
@@
$\pi\frac{d\delta r}{d\phi}=\gamma_{L}\delta r,$ (52)
its solutions are of an exponential form, $\delta r_{n}=\delta r_{0}e^{\gamma_{L}n}$, where recall that $n\equiv\phi/(2\pi)$ is the number of half-turns. In this expression $\gamma_{L}$ is the critical exponent controlling the (exponential) drift of nearly-bound orbits with $n$, and can be traded, using the critical curve relation $C^{\prime}(r_{p})A(r_{p})-C(r_{p})A^{\prime}(r_{p})=0$ [52], by another exponent $\lambda$ which captures the same drift but now in coordinate time $t$ as $\gamma_{L}=\pi b_{c}\lambda$, where the such exponent reads as [53]
$\lambda=\pi\left(\frac{1}{2}\frac{AC^{\prime\prime}-A^{\prime\prime}C}{AB}\right)^{\frac{1}{2}}.$ (53)
For the Schwarzschild geometry $\gamma_{L}=\pi$, indicating that an initial perturbation grows (in radial distance) by a factor $e^{\gamma_{L}}$ with each half-orbit around the black hole. For the Cotton geometry, the Lyapunov exponent reads
$\gamma_{L}=\pi(1+\gamma r_{p})^{\frac{1}{2}}.$ (54)
In Fig. 10 we compare the Lyapunov exponent of the CG geometry, as described by Eq.(54), against that of the Schwarzschild geometry.
Figure 10: Lyapunov exponent as a function of $\gamma$ (blue, dashed), as described by equation (54), in contrast with its Schwarzschild counterpart (red, dotted).
This result supports the qualitative analysis done in the previous sections. The $\lambda$ exponent of the CG black hole, differs from the Schwarzschild black hole by an order of magnitude proportional to $\gamma$. Thus, despite the exponential behaviour driven by the Lyapunov exponent, when $\gamma M\sim 10^{-12}$ the deviations between these two geometries are captured by photon rings of a similar order, requiring many half-orbits around the black hole before becoming significantly different. The gravitational lensing causing the $n=1$ and $n=2$ photon rings is far from being sensitive to the contribution of $\gamma$. Thus, one cannot hope to be able to distinguish between the Schwarzschild and CG geometries via their $n=1$ and $n=2$ ring features (i.e. relative intensities, relative widths), which are the ones that could be within reach of future interferometric detectors [54].
@@ VIII Conclusion
To study the impact of this result in geodesic effects, we first computed the exact expression for the deflection angle in the form of elliptic integrals. There are two important considerations about this result. Firstly, in order to obtain the result of the Schwarzschild solution, one must first assume $\gamma=0$ and additionally $u_{R}$ and $u_{S}=0$ (which corresponds to the consideration of infinite distances). This is because the result for Schwarzschild is usually determined using the definition for infinite distances, which is not possible in the case of Cotton gravity given its non-asymptotically flat character. Secondly, the deflection angle can be written as a function of the impact parameter $b$, though under an explicit cumbersome form. We furthermore computed the deflection angle following the approach of Bozza for the strong-field limit, searching for the roots of the deflection function $G(z)$ and verifying that the resulting expression smoothly recovers the right weak-field limit, where $M/b<<1$. We have considered a linear approximation for the parameter $\gamma$ in addition to the series with respect to the mass, and checked the consistence of the analytical approximation with the numerical integration of the deflection angle. Overall, we find that the deflection angle remains unchanged for values of the CG parameter $\gamma M<10^{-5}$.
@@
In looking for further settings to search for the presence of any differences, we studied the isoradial curves of a CG black hole against those of a Schwarzschild black hole, utilizing the generalization developed in [50] of the Poutanen formula of Ref. [51]. We found that there is a full overlap of the isoradial curves between the CG and Schwarzschild geometries, confirming the absence of any visual differences in this regard. We also considered the theoretical Lyapunov exponent for a Cotton gravity black hole, associated to the instability scale of photon rings. We derived its exact expression by perturbing the photon equation of motion about the photon sphere radius. The CG theoretical Lyapunov exponent was found to differ from Schwarzschild’s by a factor proportional to CG’s $\gamma$ parameter, and thus photon rings do not help in discriminating between both geometries, at least for those rings ($n=1$ and $n=2$) that could be expected to be measured in the last decades via interferometric means.
@@ Acknowledgements.
MER thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil, for partial financial support. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. FSNL acknowledges support from the Fundação para a Ciência e a Tecnologia (FCT) Scientific Employment Stimulus contract with reference CEECINST/00032/2018, and funding through the research grants UIDB/04434/2020, UIDP/04434/2020 and PTDC/FIS-AST/0054/2021. DRG is supported by the Spanish Agencia Estatal de Investigación Grant No. PID2022-138607NB-I00 and CNS2024-154444, funded by MCIN/AEI/10.13039/501100011033, FEDER, UE, and ERDF A way of making Europe.